METU Department of Mathematics Math 303 History of Math Concepts Fall 2012

Midterm 1

November 8 2012, 17:40 – 19:00

Name and the student number:

Part I

Q 1. Recognize the following numbers and determine in which civilization they were in use.

Q 2. What does mean ”sexagesimal numeral system” ? In which cases it is used nowadays ?

Q 3. What is the most famous Egyptian mathematical papyrus: who discovered it and what is the name of the scribe ? To which time (approximately) does this papyrus belong ?

Q 4. What is special about ancient Egyptian fractions ? How Egyptians would record \frac{1}{3} ?

Q 5. Which breakthrough had happened in Ancient Greek Mathematics in compare with Egyptian and Babylonian ?

Q 6. Explain the meaning of the Greek word ”arithmos” and write it in Greek letters. Write the name of each Greek letter in this word. Q 7. What was the meaning of word ”mathema” for ancient Greeks ? Write it in Greek letters. Q 8. In which century did live Thales ? Formulate some theorem which is attributed to him. Q 9. How do you understand the World Harmony Principle of Pythagoras ? Q 10. What is the concept of quadrature and whom it is attributed to ? Q 15. What was the names of the Schools founded by Plato and by Aristotle ? In which century did they live ? Q 16. Give the names of the Platonic solids and sketch at least three of them. What are the numbers of faces and vertices of these solids ? Q 17. What are the names of polygons with n corners, where n = 5, 6, 7, 8 ?